### v1

# On Oscillatory Integral Operators Satisfying the cinematic curvature condition

Abstract.
We study the oscillatory integral operators satisfying the cinematic curvature condition. First, we formulate a conjecture for this class of operators, motivated by certain necessary conditions arising from counterexamples. We then establish an estimate for these operators by combining Wolff’s two-ends reduction with refined decoupling inequalities.
1. Introduction
1.1. Statement of results
Let $n\geq 3$ and $B^{m}_{r}(y)\subset\mathbb{R}^{m}$ denote the ball of radius $r$ centered at the $y\in\mathbb{R}^{m}$.
Let
$\mathbb{A}:=\{(\xi^{\prime},\xi_{n-1})\in\mathbb{R}^{n-1}:1\leq\xi_{n-1}\leq 2,|\xi^{\prime}|\leq\xi_{n-1}\}.$
Suppose $a\in C_{c}^{\infty}(\mathbb{R}^{n}\times\mathbb{R}^{n-1})$ is non-negative and supported in $B^{n}_{1}(0)\times\mathbb{A}$. Let $\phi:B^{n}_{1}(0)\times\mathbb{A}\mapsto\mathbb{R}$ satisfy the Cinematic Curvature Condition:
$\bullet$ (H1) Homogeneous condition.
$\phi(x;\xi)$ is homogeneous of degree 1 in $\xi$ and smooth away from $\xi=0$
$\bullet$ (H2) Non-degeneracy condition.
rank $\partial^{2}_{x^{\prime}\xi}\phi(x^{\prime},x_{n};\xi)=n-1$ for all $(x^{\prime},x_{n};\xi)\in{\rm supp}\ a$ with $\xi\neq 0$.
$\bullet$ (H3) Positive definiteness condition
Consider the Gauss map $G:{\rm supp}\ a\rightarrow S^{n-1}$ by $G(x;\xi):=\frac{G_{0}(x;\xi)}{|G_{0}(x;\xi)|}$, where
$G_{0}(x;\xi):=\bigwedge_{j=1}^{n-1}\partial_{\xi_{j}}\partial_{x}\phi(x;\xi).$
Then for all $(x,\xi_{0})\in{\rm supp}\ a$,
$\partial^{2}_{\xi\xi}\langle\partial_{z}\phi(x,\xi),G(x,\xi_{0})\rangle|_{\xi=\xi_{0}}$
has rank n2n-2 with n2n-2 positive eigenvalues.
For any $\lambda\geq 1$ let $a^{\lambda}(x;\omega):=a(x/\lambda;\omega)$, $\phi^{\lambda}(x;\omega):=\lambda\phi(x/\lambda;\omega)$ and define the operator $T^{\lambda}$ by
$T^{\lambda}f(x):=\int e^{-2\pi i\phi^{\lambda}(x;\xi)}a^{\lambda}(x;\xi)f(\xi)d\xi.$
for all integrable $f\colon B^{n-1}\to\mathbb{C}$. In this case $T^{\lambda}$ is said to be an oscillatory integral operator satisfying the cinematic curvature condition (or cone condition).
The main result of this article is the following.
Theorem 1.1.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition, then the estimate
(1.1) $\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{\phi,a}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
holds uniformly for $\lambda\geq 1$ whenever
(1.2) $p>2+\frac{8}{3n-5}$
Observe that the Knapp example indicates that estimate (1.1) may fail when $n=3$ and $p\leq 4$. Consequently, the range (1.2) is sharp in the case $n=3$.
We denote by $B_{r}^{m}$ any ball of radius $r$ in $\mathbb{R}^{m}$. By an $\varepsilon$-removal trick, Theorem 1.1 can be reduced to
Proposition 1.2.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition. Then the estimate
$\|T^{\lambda}f\|_{L^{p}(B^{n}_{R})}\lesssim_{\phi,a,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
holds uniformly for $1\leq R\leq\lambda$ and all $\varepsilon>0$ whenever
(1.3) $p\geq 2+\frac{8}{3n-5}$
Although estimates for oscillatory integral operators satisfying the cinematic curvature condition may be regarded as a generalized extension of cone restriction estimates, the polynomial partitioning techniques that have recently proved effective in the study of the cone restriction problem (see [11] and [14]) do not seem to apply directly in this setting. The main obstacle lies in the fact that the transverse equidistribution estimate, which plays a crucial role in the polynomial partitioning framework, may not hold for oscillatory integral operators under cinematic curvature.
Our proof of Proposition 1.2 relies primarily on Wolff’s two-ends reduction and refined decoupling inequalities. This methodological framework has recently been applied in the study of local smoothing estimates by Gan-Wu [4] and in the the restriction problem by Wang–Wu [13]. Accordingly, our argument may be viewed as similar in spirit to their approach in the present setting of oscillatory integral operators satisfying the cinematic curvature condition.
In addition, we provide a counterexample showing that when the estimate (1.1) may fail. This example is inspired by the work of Bourgain-Guth [2] and arises from the Kakeya compression phenomenon, where tubes are arranged within a neighborhood of a low-dimensional set. A detailed discussion of this example is given in Section 2.
Theorem 1.3.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition. The estimate (1.1) may fail when
(1.4) $p<2+\frac{8}{3n-1}$ if $n$ is odd;
$p<2+\frac{8}{3n-2}$ if $n$ is even.
Estimates for oscillatory integral operators satisfying the cinematic curvature condition are closely related to several central problems in modern harmonic analysis, most notably the local smoothing problem and the cone restriction problem. These operators, by incorporating variable coefficients and geometric curvature conditions, provide a unifying analytic framework that captures many classical oscillatory phenomena while extending beyond the constant-coefficient cone scenario. Because sharp bounds in this setting would deepen our understanding of decoupling, restriction, and dispersive regularity estimates, progress on this problem has the potential to inform a broad range of techniques in Fourier analysis, partial differential equations, and geometric measure theory. Accordingly, the study of such estimates carries substantial intrinsic value, both from a conceptual standpoint and in terms of possible applications to longstanding conjectures in harmonic analysis.
1.2. Structure of the article
The structure of this paper is as follows. In Section 2 , we construct a counterexample demonstrating that the estimate (1.1) may fail when $p$ lies inside the range (1.4). Section 3 presents several preliminary reductions and methodological tools, including the narrow–broad argument and wave packet decomposition. Section 4 is devoted to the two-ends reduction. In Section 5, we complete the proof of the main theorem. Finally, Section 6 establishes an $\varepsilon$-removal lemma for oscillatory integral operators satisfying the cinematic curvature condition.
1.3. Notations
We write #E\#E for the cardinality of a finite set $E$. If $\mathcal{E}$ is a family of sets in $\mathbb{R}^{n}$, we use $\cup\mathcal{E}$ to denote $\bigcup_{E\in\mathcal{E}}E$. For $A,B\geq 0$, we write $A\lesssim B$ to mean $A\leq CB$ for some absolute constant $C$, and $A\sim B$ to mean both $A\lesssim B$ and $B\lesssim A$. For $R>1$, we write $A\lesssim_{\eta}B$ to indicate that $A\leq C_{\eta}R^{\eta}B$ for all $\eta>0$. We use $\mathrm{RapDec}(R)$ to denote a rapidly decaying quantity, meaning that for all $R\geq 1$,
$\mathrm{RapDec}(R)\leq C_{N}R^{-N}\qquad\textrm{for some very large $N$}.$
Finally, for a set $X$, we write $N_{r}(X)$ for its $r$-neighborhood.
1.4. Choice of Parameters
[…]
$R^{\delta}\ll\kappa\ll K_{\circ}\ll K\ll R\leq\lambda.$
1.5. Acknowledgments
The author would like to express sincere gratitude to Prof. Xiaochun Li for suggesting this research topic, for numerous insightful discussions, and for carefully reading earlier versions of this manuscript. His guidance and suggestions were invaluable throughout the development of this work. The author is also grateful to Prof. Shukun Wu for several helpful conversations and constructive comments that contributed to clarifying some technical aspects of this paper.
2. Necessary Conditions
In this section, we examine an example of oscillatory integral operators that satisfy the cinematic curvature condition, with the aim of demonstrating the necessity of the $p$-range stated in Theorem 1.3. The construction is motivated by the work of Bourgain–Guth [2] (see also [10] [15] [7]) and is closely related to the Kakeya compression phenomenon, in which a large collection of tubes is efficiently packed into a neighbourhood of a low-dimensional set.
Define $\bm{A}:\mathbb{R}\rightarrow\mathrm{Sym}(n-2,\mathbb{R})$ by
$\bm{A}(t):=\begin{cases}\underbrace{\begin{pmatrix}t&t^{2}\\ t^{2}&t+t^{3}\end{pmatrix}\oplus\dots\oplus\begin{pmatrix}t&t^{2}\\ t^{2}&t+t^{3}\end{pmatrix}}_{\text{$\lfloor\frac{n-2}{2}\rfloor$-fold}}\oplus\,(t),&\text{if $n$ is odd},\\[4.0pt] \underbrace{\begin{pmatrix}t&t^{2}\\ t^{2}&t+t^{3}\end{pmatrix}\oplus\dots\oplus\begin{pmatrix}t&t^{2}\\ t^{2}&t+t^{3}\end{pmatrix}}_{\text{$\lfloor\frac{n-2}{2}\rfloor$-fold}},&\text{if $n$ is even}.\end{cases}$
Let $a$ be a bump function such that $a(x,\xi)=1$ for all $(x,\xi)\in B^{n}_{1/2}(0)\times 0.8\mathbb{A}$ where
$0.8\mathbb{A}:=\{(\xi^{\prime},\xi_{n-1})\in\mathbb{R}^{n-1}:1.1\leq\xi_{n-1}\leq 1.9,|\xi^{\prime}|\leq 0.8\xi_{n-1}\}.$
and
$\phi(x^{\prime},x_{n};\xi^{\prime},\xi_{n-1}):=\langle x^{\prime},(\xi^{\prime},\xi_{n-1})\rangle+\frac{1}{2\xi_{n-1}}\langle\bm{A}(x_{n})\xi^{\prime},\xi^{\prime}\rangle$
We claim that $\phi$ satisfies (H1), (H2) and (H3).
Proposition 2.1.
$\phi$ satisfies (H1), (H2) and (H3).
Proof.
It is clear that $\phi$ satisfies (H1) and (H2). Therefore, it remains to verify that $\phi$ also satisfies (H3). For any $x=(x^{\prime},x_{n})\in B^{n}_{1}(0)$ and $\xi=(\xi^{\prime},\xi_{n-1}),\xi_{0}=(\xi_{0}^{\prime},\xi_{0,n-1})\in\mathbb{A}$,
$\partial_{x}\phi(x,\xi)=(\xi,\frac{\langle A^{\prime}(x_{n})\xi^{\prime},\xi^{\prime}\rangle}{2\xi_{n-1}}),\qquad G(x,\xi_{0})=C(x,\xi_{0})(-\frac{A^{\prime}(x_{n})\xi_{0}^{\prime}}{\xi_{0,n-1}},\frac{A(x_{n})\xi_{0}^{\prime}}{\xi^{2}_{0,n-1}},1)$
where $C(x,\xi_{0})>0$ denotes a normalization constant. Now
$\partial^{2}_{\xi\xi}\langle\partial_{\phi}(x,\xi),G(x,\xi_{0})\rangle|_{\xi=\xi_{0}}=\frac{C(x,\xi_{0})}{\xi_{0,n-1}}\begin{pmatrix}A^{\prime}(x_{n})&-\frac{A^{\prime}(x_{n})\xi_{0}^{\prime}}{\xi_{0,n-1}}\\ -(\frac{A^{\prime}(x_{n})\xi_{0}^{\prime}}{\xi_{0,n-1}})^{\top}&\frac{\langle A^{\prime}(x_{n})\xi_{0}^{\prime},\xi_{0}^{\prime}\rangle}{\xi_{0,n-1}^{2}}\end{pmatrix}$
Let
$M=\begin{pmatrix}A^{\prime}(x_{n})&-\frac{A^{\prime}(x_{n})\xi_{0}^{\prime}}{\xi_{0,n-1}}\\ -(\frac{A^{\prime}(x_{n})\xi_{0}^{\prime}}{\xi_{0,n-1}})^{\top}&\frac{\langle A^{\prime}(x_{n})\xi_{0}^{\prime},\xi_{0}^{\prime}\rangle}{\xi_{0,n-1}^{2}}\end{pmatrix}$
It remains to show that $M$ has rank n2n-2 with n2n-2 positive eigenvalues. Obviously, the rank of this matrix is n2n-2. In addition, for any $v=(v^{\prime},v_{n-1})\in\mathbb{R}^{n-1}$ with $v\neq 0$
$v^{\top}Mv=\\ \left(v^{\prime}-\frac{A^{\prime}(x_{n})\xi_{0}^{\prime}}{\xi_{0,n-1}}v_{n-1}\right)^{\top}A^{\prime}(x_{n})\left(v^{\prime}-\frac{A^{\prime}(x_{n})\xi_{0}^{\prime}}{\xi_{0,n-1}}v_{n-1}\right)\geq 0$
since $A^{\prime}(x_{n})$ is positive definite. This finishes the proof. ∎
Cover $B^{n-2}_{4/5}(0)$ by finite-overlapping caps $\theta$ of diameter $\sim\lambda^{-1/2}$. Let $\{\psi_{\theta}\}$ be a smooth partition of unity subordinate to this cover. For each $\theta$ let $\omega_{\theta}$ denote the center of $\theta$ and define $f_{\theta}:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ by
$f_{\theta}(\xi^{\prime},\xi_{n-1}):=\begin{cases}e^{2\pi i\lambda\langle v_{\theta},\xi\rangle}\,\psi_{\theta}(\xi^{\prime}/\xi_{n-1}),&\text{if }(\xi^{\prime},\xi_{n-1})\in\mathbb{A},\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
where $\xi=(\xi^{\prime},\xi_{n-1})$ and
$v_{\theta,2j-1}:=-\omega_{\theta,2j},\qquad v_{\theta,2j}=v_{\theta,n-2}=v_{\theta,n-1}=0,\qquad\mathrm{for}\quad 1\leq j\leq\left\lfloor\frac{n-2}{2}\right\rfloor$
For $(u^{\prime},u_{n-1},x_{n})\in B^{n}_{\lambda}(0)$ let
$\gamma^{\lambda}_{\theta}(u^{\prime},u_{n-1},x_{n}):=(u^{\prime}-\lambda\mathbb{A}(\frac{x_{n}}{\lambda})\omega_{\theta},u_{n-1}+\frac{\lambda}{2}\langle\mathbb{A}(\frac{x_{n}}{\lambda})\omega_{\theta},\omega_{\theta}\rangle)$
Let $\tilde{\theta}\subset\mathbb{R}^{n-1}$ be a box of dimensions $1\times\lambda^{-1/2}\times...\times\lambda^{-1/2}$ such that
$\{(\xi^{\prime},\xi_{n-1}):\frac{\xi^{\prime}}{\xi_{n-1}}\in\theta,1\leq\xi_{n-1}\leq 2\}\subset\tilde{\theta}$
and $\bar{T}_{\theta}$ denote the dual box of $\tilde{\theta}$ centered at $\lambda v_{\theta}$. Define $T_{\theta}\subset\mathbb{R}^{n}$ by
$T_{\theta}:=\{(\gamma^{\lambda}_{\theta}(u,x_{n}),x_{n}):u\in\lambda^{\delta}\bar{T}_{\theta}\ ,\quad-\lambda\leq x_{n}\leq\lambda\}\cap B^{n}_{\lambda}(0)$
For each $\theta$, there exists a significant portion of $T_{\theta}$ that guarantees $T^{\lambda}f_{\theta}$ maintains a nontrivial size.
Proposition 2.2.
For each $\theta$ there exists $S\subset T_{\theta}$ such that $|S|\gtrsim\lambda^{n/2}$ and for all $x\in S$
(2.1) $T^{\lambda}f_{\theta}(x)\gtrsim\lambda^{-O(\delta)-\frac{n-2}{2}}$
Proof.
Fix $t\in[-\lambda/2,\lambda/2]$ and define $T_{\theta,t}:=T_{\theta}\cap\{x_{n}=t\}$. It suffices to show that there exists $S_{t}\subset T_{\theta,t}$ with $|S_{t}|\gtrsim\lambda^{\frac{n-2}{2}}$ and (2.1) holds for all $x\in S_{t}$. Since
$T^{\lambda}f_{\theta}(x^{\prime},t)=\int e^{-2\pi i\langle x^{\prime},\xi\rangle}e^{-\frac{\pi i}{\xi_{n-1}}\langle\bm{A}(t/\lambda)\xi^{\prime},\xi^{\prime}\rangle}a^{\lambda}(x^{\prime},t,\xi)f_{\theta}(\xi)d\xi=\hat{f_{\theta,x^{\prime},t}}(x^{\prime})$
where
$f_{\theta,x^{\prime},t}(\xi)=e^{-\frac{\pi i}{\xi_{n-1}}\langle\bm{A}(t/\lambda)\xi^{\prime},\xi^{\prime}\rangle}a^{\lambda}(x^{\prime},t,\xi)f_{\theta}(\xi)$
Since $T^{\lambda}f_{\theta}$ is essentially supported in $T_{\theta}$ (see Lemma 3.6), the Plancherel theorem gives
(2.2) $\|T^{\lambda}f_{\theta}\|^{2}_{L^{2}(T_{\theta,t})}\sim\|f_{\theta}\|^{2}_{L^{2}(\mathbb{R}^{n-1})}\sim\lambda^{-\frac{n-2}{2}}$
Also,
(2.3) $|T^{\lambda}f_{\theta}|\leq\int f_{\theta}\leq\lambda^{-\frac{n-2}{2}}$
Let $S_{t}$ be the set of all points in $T_{\theta,t}$ satisfying (2.1). By (2.2) and (2.3),
$\lambda^{-\frac{n-2}{2}}\lesssim\|T^{\lambda}f_{\theta}\|^{2}_{L^{2}(T_{\theta,t})}\lesssim|S_{t}|\lambda^{-(n-2)}$
Thus, $|S_{t}|\gtrsim\lambda^{\frac{n-2}{2}}$. ∎
By a slight abuse of notation, we continue to use $T_{\theta}$ to denote this significant portion. Thus,
(2.4) $T^{\lambda}f_{\theta}\gtrsim\lambda^{-O(\delta)-\frac{n-2}{2}}\chi_{T_{\theta}}$
Next, we show that $\bigcup_{\theta:\lambda^{-1/2}\mathrm{-cap}}T_{\theta}$ is contained in a neighborhood of a low-dimensional set, illustrating the Kakeya compression phenomenon.
Proposition 2.3.
There exists an algebraic variety $Z$ of dimension $m:=n-\lfloor\frac{n-2}{2}\rfloor$ and degree $O_{n}(1)$ such that
(2.5) $\bigcup_{\theta:\lambda^{-1/2}\mathrm{-cap}}T_{\theta}\subset N_{\lambda^{1/2+\delta}}(Z)\cap B^{n}_{\lambda}(0)$
Proof.
Let
$c(T_{\theta}):=\{(\gamma^{\lambda}_{\theta}(\lambda v_{\theta},x_{n}),x_{n}):,\quad-\lambda\leq x_{n}\leq\lambda\}\cap B^{n}_{\lambda}(0)$
denote the center of $T_{\theta}$. A direct computation gives that $c(T_{\theta})$ lies in the common zero set $Z$ of the polynomials
$P_{j}:=\lambda x_{2j}-x_{2j-1}x_{n}\qquad\mathrm{for}\quad 1\leq j\leq\lfloor\frac{n-2}{2}\rfloor$
for all $\theta$. Since $T_{\theta}\subset N_{\lambda^{1/2+\delta}}(c(T_{\theta}))$, it follows that (2.5) holds. ∎
We now construct a function $f$ that causes estimate (1.1) to fail whenever $p$ lies in the range(1.4).
Proposition 2.4.
There exists a map $\epsilon:\{\theta\}\mapsto\{-1,1\}$ such that
$f:=\sum_{\theta:\lambda^{-1/2}\mathrm{-cap}}\epsilon(\theta)f_{\theta}$
forces estimate (1.1) to fail whenever $p$ lies within the range(1.4).
Proof.
Let $\epsilon_{\theta}$ be uniformly distributed independent random signs and
$f:=\sum_{\theta:\lambda^{-1/2}\mathrm{-cap}}\epsilon_{\theta}f_{\theta}$
For $x\in B_{\lambda}^{n}(0)$ Khintchine’s theorem (see [16], page 44) gives,
(2.6) $\mathbb{E}[|T^{\lambda}f(x)|]\sim(\sum_{\theta:\lambda^{-1/2}\mathrm{-cap}}|T^{\lambda}f_{\theta}(x)|^{2})^{1/2}\gtrsim\lambda^{-O(\delta)-\frac{n-2}{2}}(\sum_{\theta:\lambda^{-1/2}\mathrm{-cap}}\chi_{T_{\theta}})^{1/2}$
Let $p>2$. By Hölder inequality and Proposition 2.2,
(2.7) $\|(\sum_{\theta:\lambda^{-1/2}\mathrm{-cap}}\chi_{T_{\theta}})^{1/2}\|_{L^{p}(\mathbb{R}^{n})}|\bigcup_{\theta:\lambda^{-1/2}\mathrm{-cap}}T_{\theta}|^{1/2-1/p}\geq\\ \|(\sum_{\theta:\lambda^{-1/2}\mathrm{-cap}}\chi_{T_{\theta}})^{1/2}\|_{L^{2}(\mathbb{R}^{n})}\gtrsim\lambda^{\frac{n-1}{2}}$
By Proposition 2.3,
(2.8) $|\bigcup_{\theta:\lambda^{-1/2}\mathrm{-cap}}T_{\theta}|^{1/2-1/p}\lesssim\lambda^{\frac{n+m}{2}(\frac{1}{2}-\frac{1}{p})+O(\delta)}$
By Minkowski’s inequality, (2.7) and (2.8),
$\mathbb{E}[\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}]\geq\|\mathbb{E}[|T^{\lambda}f|]\|_{L^{p}(\mathbb{R}^{n})}\gtrsim\lambda^{-O(\delta)+\frac{1}{2}-\frac{n+m}{2}(\frac{1}{2}-\frac{1}{p})}$
Thus, there exists a map $\epsilon:\{\theta\}\mapsto\{-1,1\}$ such that $f:=\sum_{\theta:\lambda^{-1/2}\mathrm{-cap}}\epsilon(\theta)f_{\theta}$ satisfies
(2.9) $\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}\gtrsim\lambda^{-O(\delta)+\frac{1}{2}-\frac{n+m}{2}(\frac{1}{2}-\frac{1}{p})}$
Suppose the estimate (1.1)) holds. Since $\|f\|_{L^{p}(\mathbb{R}^{n-1})}\sim 1$, (2.9) gives
$\lambda^{-O(\delta)+\frac{1}{2}-\frac{n+m}{2}(\frac{1}{2}-\frac{1}{p})}\lesssim 1$
which is equivalent to
$p\geq 2+\frac{4}{n+m-2}$
This gives Theorem 1.3. Observe that when $n\leq 5$, the Knapp example yields a larger range of exponents for which estimate (1.1) may fail. Consequently, the necessary conditions established in this paper are not optimal. It is reasonable to anticipate that a sharper counterexample will be identified in subsequent work.
3. Preliminaries
In this section, we introduce the notation and basic properties for oscillatory integral operators satisfying the cinematic curvature condition. Throughout the remainder of the paper, we write $x=(x^{\prime},x_{n})\in\mathbb{R}^{n}$ and $\xi=(\xi^{\prime},\xi_{n-1})\in\mathbb{R}^{n-1}$.
3.1. Some Basic Reductions
One can reduce Proposition 1.2 to the case $1\leq R\leq\lambda^{1-\varepsilon}$ and $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$. Moreover, after an appropriate rescaling, we may further assume that $B^{n}_{R}=B^{n}_{R}(0)$ and
(3.1) $\|\partial_{x,\xi}^{\beta}a\|_{L^{\infty}(B^{n}_{1}(0)\times\mathbb{A})}\lesssim_{\beta}1\qquad\textrm{for all $1\leq|\beta|\leq N$.}$
where $N=N_{\varepsilon,M,p}\in\mathbb{N}$ is some large number depending only on the dimension $n$ and the fixed choice of $\varepsilon,M$ and $p$. Hence, to prove Proposition 1.2,it suffices to show that:
Proposition 3.1.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition. Then the estimate
$\|T^{\lambda}f\|_{L^{p}(B^{n}_{R}(0))}\lesssim_{\phi,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{p}(B^{n-1})}$
holds uniformly for $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$, $1\leq R\leq\lambda^{1-\varepsilon}$ and $\varepsilon>0$, provided that the amplitude $a$ satisfies (3.1) and the exponent $p$ lies in the range (1.3).
To obtain uniform estimates for a family of oscillatory integral operators satisfying the cone condition, it is necessary to strengthen Proposition 3.1 This strengthening is essential for implementing an induction-on-scales argument, since a single operator $T^{\lambda}$ is not stable under rescaling. Nevertheless, the rescaling procedure preserves certain structural features of this class of operators, allowing one to carry out the induction simultaneously over a larger family. In what follows, we work within the framework introduced in [1, Section 2.3].
Fix a small constant $c_{\text{par}}>0$ be a small fixed constant. For ${\bf A}=(A_{1},A_{2},A_{3})\in[1,\infty)^{3}$, we impose the following conditions on the phase function:
  1. $(\text{H1}_{{\bf A}})$
    $\|\partial^{2}_{\xi x^{\prime}}\phi(x;\xi)-I_{n-1}\|_{\mathrm{op}}\leq c_{\text{par}}A_{1}$ for all $(x;\xi)\in B^{n}_{1}(0)\times\mathbb{A}$.
  2. $(\text{H2}_{{\bf A}})$
    $\|\partial^{2}_{\xi^{\prime}\xi^{\prime}}\partial_{x_{n}}\phi(x;\xi)-\frac{1}{\xi_{n-1}}I_{n-2}\|_{\mathrm{op}}\leq c_{\text{par}}A_{2}$ for all $(x;\xi)\in B^{n}_{1}(0)\times\mathbb{A}$.
  3. $(\textup{D1}_{\bf A})$
    $\|\partial_{\xi}^{\beta}\partial_{x_{k}}\phi\|_{L^{\infty}(B_{1}^{n}(0)\times\mathbb{A})}\leq c_{\text{par}}A_{1}$ for all $1\leq k\leq n-1$ and $\beta\in\mathbb{N}_{0}^{n-1}$ with $2\leq|\beta|\leq 3$ satisfying $|\beta^{\prime}|\geq 2$;
    $\|\partial_{\xi^{\prime}}^{\beta^{\prime}}\partial_{x_{k}}\phi\|_{L^{\infty}(B_{1}^{n}(0)\times\mathbb{A})}\leq\frac{c_{\text{par}}}{2n}A_{1}$ for all $1\leq k\leq n-1$ and $\beta^{\prime}\in\mathbb{N}_{0}^{n-2}$ with $|\beta^{\prime}|=3$.
  4. $(\textup{D2}_{\bf A})$
    $\|\partial_{\xi}^{\beta}\partial_{z}^{\alpha}\phi\|_{L^{\infty}(B_{1}^{n}(0)\times\mathbb{A})}\leq\frac{c_{\text{par}}}{2n}A_{3}$ for all $(\alpha,\beta)\in\mathbb{N}_{0}^{n}\times\mathbb{N}_{0}^{n-1}$ with $2\leq|\alpha|\leq 4N$ and $1\leq|\beta|\leq 4N+2$ satisfying $1\leq|\beta|\leq 4N$ or $|\beta^{\prime}|\geq 2$.
  5. $(\textup{M}_{\bf A})$
    ${\rm dist}({\rm supp}_{x}a,\mathbb{R}^{n}\setminus B_{1}^{n})\geq\frac{1}{4A_{3}}$.
A pair $(\phi,a)$ is said to be of type $\bf A$} if it satisfies the conditions $(\textup{H1}_{\bf A}),(\textup{H2}_{\bf A}),\\ (\textup{D1}_{\bf A}),(\textup{D2}_{\bf A})$ and $(\textup{M}_{\bf A})$ (in addition to (H1) and (H2)) is said to be of type $\bf A$}. By decomposing the amplitude function and applying suitable rescaling arguments, one may reduce to the normalized case ${\bf A}=\mathbf{1}:=(1,1,1)$; see [1, Section 2.5] for details.
In the remainder of this paper, we assume that the operator $T^{\lambda}$ is of type 1. Under this normalization, Proposition 3.1 admits the following strengthened formulation.
Proposition 3.2.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
holds uniformly for $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$, $1\leq R\leq\lambda^{1-\varepsilon}$ and $\varepsilon>0$.
3.2. Wave packet decomposition
Fix $0<\varepsilon\ll 1$. In this subsection, we recall the wave packet decomposition for a function $f\in\mathcal{S}(\mathbb{R}^{n-1})$ at scale $R$ satisfying $1\ll R\leq\lambda^{1-\varepsilon}$, following the presentation in [4, Section 3.3].
Condition (H1) guarantees that the mapping $\partial_{\xi}\phi(\cdot,x_{n};\xi)$ is locally invertible. More precisely, for each fixed pair $(x_{n},\xi)$, the map $x^{\prime}\mapsto\partial_{\xi}\phi(x;\xi)$ has a non-vanishing Jacobian determinant. After a suitable rescaling, we may further assume that
(3.2) $\partial_{\xi}\phi(\cdot,x_{n};\xi)\quad\text{defines a diffeomorphism on}\quad B_{1}^{n-1}(0)\quad\text{for all}(x_{n},\xi)\in[-1,1]\times\mathbb{A}.$
Consequently, this mapping admits a well-defined inverse, which we introduce below.
Definition 3.3.
[4, Definition 3.3] For $(x_{n},\xi)\in[-1,1]\times\mathbb{A}$, let $\Phi(u,x_{n};\xi)$ be the unique solution to
(3.3) $\partial_{\xi}\phi(\Phi(u,x_{n};\xi),x_{n};\xi)=u.$
Here, we may assume the equation holds for $u$ in a small neighborhood of the origin.
Since the phase function $\phi$ is homogeneous of degree one in $\xi$, we have $\partial_{\xi}\phi(u;\xi)=\partial_{\xi}\phi(u;s\xi)$ for any $s>0$. Consequently,
(3.4) $\Phi(u;\xi)=\Phi(u;s\xi)$
Moreover, by invoking the inverse function theorem together with conditions $(\textup{D1}_{\bf A})$ and $(\textup{D2}_{\bf A})$, we obtain uniform $L^{\infty}$ bounds for $\Phi$ and its derivatives. Differentiating (3.3) with respect to $u$ yields
$\partial^{2}_{\xi,x^{\prime}=\Phi(u,x_{n};\xi)}\phi\cdot\partial_{u}\Phi=I_{n}.$
As a consequence,
$\det\partial_{u}\Phi(u,x_{n};\xi)\gtrsim 1,$
That is, the matrix $\partial_{u}\Phi(u,x_{n};\xi)$ is quantitatively non-singular.
Define $\gamma^{\lambda}:\mathbb{R}^{n-1}\times\mathbb{R}\times\mathbb{A}\mapsto\mathbb{R}^{n-1}$ by
$\gamma^{\lambda}(u,x_{n};\xi):=\lambda\Phi(u/\lambda,x_{n}/\lambda;\xi).$
Then one has
(3.5) $\partial_{\xi}\phi^{\lambda}(\gamma^{\lambda}(u,x_{n};\xi),t;\xi)=u.$
For each fixed pair $(u,\xi)$, the mapping $x_{n}\mapsto(\gamma^{\lambda}(u,x_{n};\xi),x_{n})$, with $x_{n}\in[-R,R]$, traces out a curve. The wave packets considered here will be defined as certain non-isotropic neighborhoods surrounding these curves.
We proceed to describe the wave packet decomposition of $T^{\lambda}f$ at scale $R$. Cover $B^{n-2}_{1}(0)$ by finite-overlapping caps $\theta$ of diameter $\sim R^{-1/2}$. Let $\Theta_{R^{-1/2}}:=\{\theta\}$ and $\{\psi_{\theta}\}$ be a smooth partition of unity subordinate to this cover. For each $\theta$ let $\omega_{\theta}$ denote the center of $\theta$ and define $\tilde{\psi_{\theta}}$ by
$\tilde{\psi_{\theta}}(\xi^{\prime},\xi_{n-1}):=\begin{cases}\psi_{\theta}(\xi^{\prime}/\xi_{n-1}),&\text{if }(\xi^{\prime},\xi_{n-1})\in\mathbb{A},\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
Let $f_{\theta}:=\tilde{\psi}_{\theta}f$. Now we can write
$f=\sum_{\theta\in\Theta_{r^{-1/2}}}f_{\theta}$
Let $\tilde{\theta}\subset\mathbb{R}^{n-1}$ be a box of dimensions $1\times R^{-1/2}\times...\times R^{-1/2}$ such that
$\{(\xi^{\prime},\xi_{n-1}):\frac{\xi^{\prime}}{\xi_{n-1}}\in\theta,1\leq\xi_{n-1}\leq 2\}\subset\tilde{\theta}$
In addition, we denote $C\tilde{\theta}$ by a $1\times CR^{-1/2}\times...\times CR^{-1/2}$ box satisfying
$\{(\xi^{\prime},\xi_{n-1}):\frac{\xi^{\prime}}{\xi_{n-1}}\in C\theta,0.9\leq\xi_{n-1}\leq 2.1\}\subset C\tilde{\theta}$
We first introduce some notations before decomposing in physical space.
Definition 3.4.
[4, Definition 3.4] Let $U\subset\mathbb{R}^{m}$ be a box of dimensions $a_{1}\times\dots\times a_{m}$. Define $U^{*}$ to be the dual box of $U$ centered at the origin of dimensions $a_{1}^{-1}\times\dots\times a_{m}^{-1}$, and with edges parallel to the corresponding edges of $U$.
Definition 3.5.
[4, Remark 3.6] For a slab $U\subset\mathbb{R}^{n-1}$ of dimensions $Rs^{2}\times Rs\times\dots\times Rs$ with $s\in[R^{-1/2},1]$ and a constant $C>0$, define $CU$, the non-isotropic $C$-dilation of $U$, to be a slab of dimensions $C^{2}Rs^{2}\times CRs\times\dots\times CRs$ with the same center as $U$.
For two slabs $U_{1},U_{2}\subset\mathbb{R}^{n}$, we say they are comparable if
$\frac{1}{C}U_{1}\subset U_{2}\subset CU_{1}.$
Here, $C=C_{\phi}$ is a constant that may vary from line to line, but eventually only depends on $\phi$.
We now carry out the wave packet decomposition of $f$. Fix $\theta\in\Theta_{R^{-1/2}}$. We cover $\mathbb{R}^{n-1}$ by a finitely overlapping family of $R^{2\delta}\times R^{1/2+\delta}\times\cdots\times R^{1/2+\delta}$ slabs that are parallel to $R^{\delta}\tilde{\theta}^{*}$. Denote this collection by
(3.6) $\mathcal{T}_{\theta}^{\flat}:=\{T^{\flat}\}.$
Let $\{\eta_{T^{\flat}}\}$ be a smooth partition of unity subordinate to this cover. Then we may decompose $\widehat{f}$ as
$\widehat{f}=\sum_{\theta\in\Theta_{R^{-1/2}}}\sum_{T^{\flat}\in\mathcal{T}_{\theta}^{\flat}}\eta_{T^{\flat}}\,\widehat{f}_{\theta}.$
For each $\theta\in\Theta_{R^{-1/2}}$ and $T^{\flat}\in\mathcal{T}_{\theta}^{\flat}$, the function $\widecheck{\eta_{T^{\flat}}\widehat{f}_{\theta}}=\widecheck{\eta_{T^{\flat}}}*f_{\theta}$ is essentially supported in $2\tilde{\theta}$. Consequently,
(3.7) $f=\sum_{\theta\in\Theta_{R^{-1/2}}}\sum_{T^{\flat}\in\mathcal{T}_{\theta}^{\flat}}(\widecheck{\eta_{T^{\flat}}}*f_{\theta})\,\psi_{\tilde{\theta}}+\mathrm{Rap}(r)\|f\|_{2},$
where $\psi_{\tilde{\theta}}$ is a smooth bump function satisfying ${\rm supp}\psi_{\tilde{\theta}}\subset\tilde{3\theta}$ and $\psi_{\tilde{\theta}}(\xi)=1$ for all $\xi\in\tilde{2\theta}$.
For each $\theta\in\Theta_{R^{-1/2}}$ and $T^{\flat}\in\mathcal{T}_{\theta}^{\flat}$, we associate a tube
$T:=\bigl\{\bigl(\gamma^{\lambda}(u,x_{n};\omega_{\theta},1),x_{n}\bigr):u\in R^{\delta}T^{\flat},\ -R\leq x_{n}\leq R\bigr\}.$
and define
(3.8) $f_{T}:=(\widecheck{\eta_{T^{\flat}}}*f_{\theta})\,\psi_{\tilde{\theta}}$.
$T^{\lambda}f_{T}$ is concentrated on $T$. A more precise statement is as follows:
Lemma 3.6.
[4, Lemma 3.12] $(T^{\lambda}f_{T}){\bf{1}}_{B_{R}^{n}(0)}$ is essentially supported in $T$. In other words,
$(T^{\lambda}f_{T}){\bf{1}}_{B_{R}^{n}(0)\setminus T}={\rm RapDec}(R)\|f\|_{2}.$
With this notation (3.8), (3.7) may be rewritten as
$f=\sum_{\theta\in\Theta_{R^{-1/2}}}\sum_{T\in\mathcal{T}_{\theta}}f_{T}+\mathrm{Rap}(R)\|f\|_{2}.$
We write $\mathcal{T}_{\theta}:=\{T:T^{\flat}\in\mathcal{T}_{\theta}^{\flat}\}$ and $\mathcal{T}:=\bigcup_{\theta\in\Theta_{R^{-1/2}}}\mathcal{T}_{\theta}$.
This yields the wave packet decomposition of $f$ at scale $R$. From Plancherel theorem, we see that
Lemma 3.7.
Suppose $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$ and $\mathcal{T}^{\prime}\subset\mathcal{T}$, then
$\|\sum_{T\in\mathcal{T}^{\prime}}f_{T}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}\sim\sum_{T\in\mathcal{T}^{\prime}}\|f_{T}\|_{2}^{2}$
For each tube $T\in\mathcal{T}$ and each $R\leq t\leq R$, the cross-section $T\cap\{x_{n}=t\}$ is a rectangular box with dimensions $R^{2\delta}\times R^{1/2+\delta}\times\cdots\times R^{1/2+\delta}$. To state this more precisely, we introduce the following definition.
Definition 3.8.
[4, Definition 3.5] Given $\xi_{0}\in\mathbb{A}$ and a box $V\subset B_{R}^{n-1}(0)$, define
(3.9) $\Gamma_{V}(\xi_{0};R):=\{(\gamma^{\lambda}(u,x_{n};\xi_{0}),t):u\in V,|x_{n}|\leq R\}.$
$\Gamma_{V}(\xi_{0};R)$ is called a curved plank with base $V$ and direction $\xi_{0}$.
Using the notation of curved planks, each tube $T\in\mathcal{T}$ can be represented as
$T=\Gamma_{R^{\delta}T^{\flat}}(\omega_{\theta},1;R),$
that is, a curved plank with base $R^{\delta}T^{\flat}$ and direction $(\omega_{\theta},1)$. We refer to each such $T$ as an $R$-plank. Formally,
Lemma 3.9.
[4, Lemma 3.9] Let $s\in[R^{-1/2},1]$ and $\xi_{0}\in\mathbb{A}$. Let $V\subset\mathbb{R}^{n-1}$ be a slab of dimensions $Rs^{2}\times Rs\times\dots\times Rs$. Then for $|t|\leq R$, each $t$-slice $\Gamma_{V}(\xi_{0};R)\cap\{x_{n}=t\}$ is comparable to a $Rs^{2}\times Rs\times\dots\times Rs$-slab.
3.3. Some simple estimates
The following standard estimates will play an important role in the decoupling argument that follows. Roughly they assert that $T^{\lambda}f_{T}$ behaves essentially like a constant on $T$.
Lemma 3.10.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1). Suppose $q\geq 2$ and $g\in\mathcal{S}(\mathbb{R}^{n-1})$ is supposed in $\mathbb{A}$, then the estimate
$\|T^{\lambda}g\|_{L^{q}(\mathbb{R}^{n-1}\times\{t\})}^{q}\lesssim_{\phi,q}|{\rm supp}\ g|^{\frac{q}{2}-1}\|g\|_{2}^{q}$
holds uniformly for all $t\in\mathbb{R}$.
Two direct corollaries are
Corollary 3.11.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Suppose $a$ satisfies (3.1)). Let $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$ and $T\in\mathcal{T}$. Let $q=2+\frac{4}{n-2}$, then we have
$\|T^{\lambda}f_{T}\|_{L^{q}(B_{2R}^{n}(0))}\lesssim_{\phi}\|f_{T}\|_{L^{2}(\mathbb{R}^{n-1})}$
Proof of Corollary 3.11 assuming Lemma 3.10.
Since $f_{T}$ is supported in $\tilde{\theta}$ for some $\theta\in\Theta_{R^{-1/2}}$, we have $|{\rm supp}\ f_{T}|\lesssim R^{-\frac{n-2}{2}}$. Thus, Lemma 3.10 gives
$\|T^{\lambda}f_{T}\|_{L^{q}(B_{R}^{n}(0))}^{q}\leq\int_{-2R}^{2R}\|T^{\lambda}f_{T}\|_{L^{q}(\mathbb{R}^{n-1}\times\{t\})}^{q}dt\lesssim_{\phi}\|f_{T}\|_{L^{2}(\mathbb{R}^{n-1})}^{q}$
Corollary 3.12.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), then the estimate
$\|T^{\lambda}f\|_{L^{2}(Q)}^{2}\lesssim_{\phi}R^{1/2}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{2}$
holds uniformly for all $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$ and all cubes $Q$ of sidelength $R^{1/2}$.
To prove Lemma 3.10, we will require the following result, due to Hörmander, which generalizes the Hausdorff–Young inequality.
Lemma 3.13.
[9, Theorem 1.1] Suppose $q\geq 2$. Let $S^{\lambda}$ be an oscillatory integral operator defined as follows:
$S^{\lambda}h(x):=\int e^{-2\pi\lambda i\bar{\phi}(x,y)}\bar{a}(x,y)h(y)dy$
where $\bar{\phi}:B^{n-1}_{1}(0)\times\mathbb{A}\mapsto\mathbb{R}$ and $\bar{a}:B^{n-1}_{1}(0)\times\mathbb{A}\mapsto\mathbb{R}^{+}$ satisfy
(3.10) $\det\partial^{2}_{x,y}\bar{\phi}(x,y)>0\qquad\text{for all $(x,y)\in B^{n-1}_{1}(0)\times\mathbb{A}$}$
(3.11) $\|\partial_{x,y}^{\beta}\bar{a}\|_{L^{\infty}(B^{n-1}_{1}(0)\times\mathbb{A})}\lesssim_{\beta}1\qquad\qquad\text{for all $1\leq|\beta|\leq N-1$.}$
then we have
$\|S^{\lambda}g\|_{L^{q}(\mathbb{R}^{n-1})}\lesssim_{\phi,q}\lambda^{-\frac{n-1}{q}}\|g\|_{L^{q^{\prime}}(\mathbb{R}^{n-1})}$
where $q^{\prime}$ is the Hölder conjugate of $q$.
Proof of Lemma 3.10.
Define $S^{\lambda}g$ by
$S^{\lambda}g(u):=T^{\lambda}g(\lambda u,t)$
Since $T^{\lambda}$ satisfies (H1) and assumption (3.1), conditions (3.10) and (3.11) are satisfied. Consequently, Lemma 3.13 yields
$\|T^{\lambda}g\|_{L^{q}(\mathbb{R}^{n-1}\times\{t\})}^{q}=\lambda^{n-1}\|S^{\lambda}g\|_{L^{q}(\mathbb{R}^{n-1})}\lesssim_{\phi,q}\|g\|_{L^{q^{\prime}}(\mathbb{R}^{n-1})}$
The proof is then completed by an application of Hölder’s inequality. ∎
3.4. Mixed Norm
Since the $L^{2}$ behavior of wave packets is significantly simpler to analyze, one can reduce Proposition 3.2 to the following via a real interpolation of strict-type. Similar reductions can be found in [5, 14].
Proposition 3.14.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1) and (3.2),then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{p}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{p}}$
$\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}:=\left(\frac{1}{|3\tilde{\theta}|}\int_{3\tilde{\theta}}|f|^{2}\right)^{1/2}$
holds uniformly for $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$, $1\leq R\leq\lambda^{1-\varepsilon}$ and $\varepsilon>0$.
We now present two simple but useful results concerning $\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}$. The first is as follows:
Lemma 3.15.
Let $f\in\mathcal{S}(\mathbb{R}^{n-1})$ be supported in $\mathbb{A}$. Suppose there exists a nonempty collection $\mathbb{T}\subset\mathcal{T}$ such that
[…]
then we have
$\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}\gtrsim u^{2}R^{\frac{n-2}{4}}|\mathbb{T}|^{1/2}$
Proof.
Define $\theta^{\prime}:=\operatorname*{arg\,max}_{\theta\in\Theta_{R^{=1/2}}}|\mathbb{T}\cap\mathcal{T}_{\theta}|$. Since $\mathbb{T}$ is nonempty we have
(3.12) $|\mathbb{T}\cap\mathcal{T}_{\theta^{\prime}}|\gtrsim\max\{1,|\mathbb{T}|R^{-\frac{n-2}{2}}\}\geq\ |\mathbb{T}|^{1/2}R^{-\frac{n-2}{4}}$
Because $\Theta_{R^{-1/2}}$ is finite-overlapping and $f_{\theta}$ is supported in $3\tilde{\theta}$, Plancherel theorem gives
$\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}\geq\max_{\theta\in\Theta_{R^{-1/2}}}\|f_{\theta}\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}\gtrsim R^{\frac{n-2}{2}}\|f_{\theta^{\prime}}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}\gtrsim\\ u^{2}R^{\frac{n-2}{2}}|\mathbb{T}\cap\mathcal{T}_{\theta^{\prime}}|+{\rm RapDec}(R)\|f\|_{2}^{2}$
We can now conclude the proof combining this and (3.12). ∎
The second one is useful in the induction-on-scale arguments.
Lemma 3.16.
Let $f\in\mathcal{S}(\mathbb{R}^{n-1})$ be supported in $\mathbb{A}$. Suppose $\mathbb{T}\subset\mathcal{T}$ and let $g:=\sum_{T\in\mathbb{T}}f_{T}$, then
$\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}\lesssim\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}$
Proof.
Fix $\theta\in\Theta_{R^{-1/2}}$. Now Lemma 3.7 gives
(3.13) $\|g\|_{L^{2}(3\tilde{\theta})}^{2}=\|\sum_{T\in\mathcal{T},3\theta(T)\cap 3\theta\neq\emptyset}f_{T}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}\lesssim\\ \|\sum_{3\theta^{\prime}\cap 3\theta\neq\emptyset}f_{\theta^{\prime}}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}$
Since $\Theta_{R^{-1/2}}$ is finite-overlapping and $f_{\theta^{\prime}}$ is supported in $3\tilde{\theta^{\prime}}$, we have
$\|\sum_{3\theta^{\prime}\cap 3\theta\neq\emptyset}f_{\theta^{\prime}}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}\lesssim\|f\|_{L^{2}(9\tilde{\theta})}^{2}$
By pigeonholing, there exists $\theta^{\prime}\in\Theta_{R^{-1/2}}$ such that
$\|\sum_{3\theta^{\prime}\cap 3\theta\neq\emptyset}f_{\theta^{\prime}}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}\lesssim\|f\|_{L^{2}(3\tilde{\theta^{\prime}})}^{2}$
Combining this with (3.13)), we complete the proof.. ∎
3.5. Lorentz rescaling
Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1) and (3.2). Suppose $g\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}g\subset\tau$ for some $\tau\in\Theta_{K^{-1}}$. To estimate $\|T^{\lambda}g\|_{L^{p}(B_{R}^{n})}$ via induction on scales, we perform a rescaling under which $T^{\lambda}g$ is transformed into $\widetilde{T}^{\lambda/K^{2}}\widetilde{g}$. Here $\widetilde{T}^{\lambda/K^{2}}$ is an oscillatory integral operator satisfying the cinematic curvature condition, and $\widetilde{g}$ is a function supported in $\mathbb{A}$. Under this rescaling, the scale of the integration domain is reduced from $R$ to $R/K^{2}$, and thus we can apply
We now state the main result of this subsection in the form of a lemma. The proof follows the framework developed in [1, Section 2.5] and [4, Section 5.1].
Lemma 3.17.
Let $n\geq 3$. Suppose there exists a constant $C>0$ such that, for every function $\widetilde{g}$ with ${\rm supp}\ \widetilde{g}\subset\mathbb{A}$ and every oscillatory integral operator $T^{\lambda/K^{2}}$ satisfying the cinematic curvature condition of type $\mathbf{1}$, the estimate
(3.14) $\|T^{\lambda/K^{2}}\widetilde{g}\|_{L^{p}(B_{R/K^{2}}^{n})}^{p}\leq C\,\|\widetilde{g}\|_{2}^{2}\max_{\theta\in\Theta_{KR^{-1/2}}}\|\widetilde{g}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}+{\rm RapDec}(R/K^{2})\,\|\widetilde{g}\|_{2}^{p}$
holds whenever $a$ satisfies (3.1).
Then, for every function $g\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}\ g\subset\tau\in\Theta_{K^{-1}}$, and every oscillatory integral operator $T^{\lambda}$ satisfying the cinematic curvature condition of type $\mathbf{1}$, the estimate
$\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}C\,K^{\,2-(n-2)(p-2)}\,\|g\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}+{\rm RapDec}(R)\,\|g\|_{2}^{p}.$
holds whenever (3.1) and (3.2) is satisfied.
Proof.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1) and (3.2).
Fix $\tau\in\Theta_{K^{-1}}$. Let $V_{\tau,R}\subset B_{R}^{n}$ be an isotropic $RK^{-2}$-dilate of $\widetilde{\tau}$; thus $V_{\tau,R}$ is a box of dimensions $RK^{-2}\times RK^{-1}\times\cdots\times RK^{-1}$. We tile $\mathbb{R}^{n-1}$ by finitely overlapping translates of $V_{\tau,R}$, and denote the resulting collection by $\mathcal{B}_{\tau}^{\flat}=\{\Box^{\flat}\}$. For each $\Box^{\flat}\in\mathcal{B}_{\tau}^{\flat}$, we define a curved plank
$\Box:=\Gamma_{\Box^{\flat}}(\xi_{\tau};R)=\{(\gamma^{\lambda}(v,t;\xi_{\tau}),t;\xi_{\tau}):v\in\Box^{\flat},|t|\leq R\}.$
Let $\mathcal{B}_{\tau}:=\{\Box:\Box^{\flat}\in\mathcal{B}_{\tau}^{\flat}\}$. Observe that $\mathcal{B}_{\tau}$ forms a covering of $\mathbb{R}^{n}$, and that each
$T\in\bigcup_{\theta\in\Theta_{R^{-1/2}},\,2\theta\cap\tau\neq\emptyset}\mathcal{T}_{\theta}$
is contained in $O(1)$ many boxes $\Box\in\mathcal{B}_{\tau}$. We assign each such $T$ to a single box $\Box$ containing it, and denote by $\mathcal{T}_{\Box}$ the collection of wave packets assigned to $\Box$. This procedure yields a partition:
$\bigcup_{\theta\in\Theta_{R^{-1/2}},2\theta\cap\tau\neq\emptyset}\mathcal{T}_{\theta}=\bigcup_{\Box}\mathcal{T}_{\Box}.$
Let $g\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}\ g\subset\tau$, then
$T^{\lambda}g=\sum_{\theta\in\Theta_{R^{-1/2}},2\theta\cap\tau\neq\emptyset}T^{\lambda}g_{\theta}=\sum_{\Box}T^{\lambda}g_{\Box}+{\rm RapDec}(R)\,\|g\|_{2}$
where
$g_{\Box}:=\sum_{T\in\mathcal{T}_{\Box}}g_{T}$
Since the boxes in $\mathbb{B}_{\tau}$have finite overlap, we have
(3.15) $\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim\sum_{\Box}\|T^{\lambda}g_{\Box}\|_{L^{p}(\Box)}^{p}+{\rm RapDec}(R)\,\|g\|_{2}^{p}.$
We now apply Lorentz rescaling to each $g_{\Box}$. Define $\tilde{g_{\Box}}$ by
(3.16) $\tilde{g_{\Box}}(\eta):=K^{-(n-2)}g_{\Box}(\eta_{n-1}\omega_{\tau}+K^{-1}\eta^{\prime},\eta_{n-1})$
As demonstrated in the proof of Lemma 2.3 in [1], we have
(3.17) $T^{\lambda}g_{\Box}\circ\Upsilon_{\omega_{\tau}}^{\lambda}\circ D_{K}=\widetilde{T}^{\lambda/K^{2}}\tilde{g_{\Box}}$
where $\Upsilon^{\lambda}_{\omega_{\tau}}(u,x_{n}):=(\lambda\gamma(u/\lambda,x_{n}/\lambda;\omega_{\tau},1),x_{n})$$D_{K}(x^{\prime\prime},x_{n-1},x_{n}):=(Kx^{\prime\prime},x_{n-1},K^{2}x_{n})$ and
$\widetilde{T}^{\lambda/K^{2}}\tilde{g_{\Box}}(x):=\int_{\mathbb{R}^{n}}e^{i\widetilde{\phi}^{\lambda/K^{2}}(x;\eta)}\widetilde{a}^{\lambda}(x;\eta)\tilde{g_{\Box}}(\eta)\mathrm{d}\eta.$
Here the phase function $\widetilde{\phi}$ and the amplitude function $\widetilde{a}(x;\eta)$ are given by
$\widetilde{\phi}(x;\eta):=\langle x^{\prime},\eta\rangle+\int_{0}^{1}(1-r)\langle\partial^{2}_{\xi^{\prime}\xi^{\prime}}\phi(\Upsilon_{\omega_{\tau}}(D^{\prime}_{K^{-1}}x^{\prime},x_{n});\eta_{n-1}\omega_{\tau}+rK^{-1}\eta^{\prime},\eta_{n-1})\eta^{\prime},\eta^{\prime}\rangle\mathrm{d}r$
$\widetilde{a}(x;\eta):=a(\Upsilon_{\omega_{\tau}}(D^{\prime}_{K^{-1}}x^{\prime};x_{n});\eta_{n-1}\omega_{\tau}+K^{-1}\eta^{\prime},\eta_{n-1})$
where
$D^{\prime}_{K^{-1}}(x^{\prime\prime},x_{n-1}):=(K^{-1}x^{\prime\prime},K^{-2}x_{n-1})$
$\Upsilon_{\omega_{\tau}}(u,x_{n}):=(\gamma(u,x_{n};\omega_{\tau},1),x_{n})$
For each $\Box\in\mathcal{B}_{\tau}$, $(\Upsilon^{\lambda}_{\omega_{\tau}}\circ D_{K})^{-1}(\Box)$ is a $R/K^{2}$-ball. We denote this ball with $B^{n}_{R/K^{2},\Box}$. By (3.17), we can now rewrite (3.15) as
(3.18) $\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim K^{n}\sum_{\Box}\|\widetilde{T}^{\lambda/K^{2}}\tilde{g_{\Box}}\|_{L^{p}(B^{n}_{R/K^{2},\Box})}^{p}+{\rm RapDec}(R)\,\|g\|_{2}^{p}.$
[…]
(3.19) $\|\widetilde{T}^{\lambda/K^{2}}\tilde{g_{\Box}}\|_{L^{p}(B^{n}_{R/K^{2},\Box})}^{p}\lesssim_{\phi}C\,\|\tilde{g_{\Box}}\|_{2}^{2}\max_{\theta\in\Theta_{KR^{-1/2}}}\|\tilde{g_{\Box}}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}\\ +{\rm RapDec}(R/K^{2})\,\|\tilde{g_{\Box}}\|_{2}^{p}$
By (3.16) and Lemma 3.16,
(3.20) $\|\tilde{g_{\Box}}\|_{2}^{2}=K^{-(n-2)}\|g_{\Box}\|_{2}^{2}$
(3.21) $\max_{\theta\in\Theta_{KR^{-1/2}}}\|\tilde{g_{\Box}}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{2}\leq K^{-2(n-2)}\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{2}$
Combine (3.18), (3.19), (3.20) and (3.21),
$\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}K^{2-(n-2)(p-2)}\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\sum_{\Box}\|g_{\Box}\|_{2}^{2}+{\rm RapDec}(R)\,\|g\|_{2}^{p}.$
Invoking Lemma 3.7, we conclude the proof. ∎
3.6. Simplification of phase function
From [4, Proof of Lemma 5.4], one can reduce Proposition 3.14 to the case
(3.22) $\phi(x;\xi)=\langle x^{\prime},\xi\rangle+x_{n}\frac{h(\xi^{\prime})}{\xi_{n-1}}+\mathcal{E}(x;\xi^{\prime}/\xi_{n-1})\xi_{n-1}$
where $h$ is quadratic and the error term is bounded by
(3.23) $\mathcal{E}(x;\omega)\lesssim|x_{n}||\omega|^{3}+|x|^{2}|\omega|^{2}$
Formally,
Proposition 3.18.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), (3.2), (3.22) and (3.23), then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{p}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{p}}$
$\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}:=\left(\frac{1}{|3\tilde{\theta}|}\int_{3\tilde{\theta}}|f|^{2}\right)^{1/2}$
holds uniformly for $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$, $1\leq R\leq\lambda^{1-\varepsilon}$ and $\varepsilon>0$.
3.7. Broad Norm
We recall the definition of the broad norm, which plays a central role in the subsequent decoupling analysis in this article. Related definitions can be found in [2, 7, 4, 5, 11].
Let $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), (3.2), (3.22) and (3.23). Suppose $p\geq 2$.
Fix $A\geq 1$ and cover $B_{\lambda}^{n}(0)$ by a finitely overlapping collection $\mathcal{B}$ of balls of radius $K^{2}$. For each ball $B_{K^{2}}^{n}(y)\in\mathcal{B}$ with $y\in B^{\lambda}(0)$, define
(3.24) $\mu(B_{K^{2}}^{n}(y)):=\min_{v_{1},\dots,v_{A}\in S^{n-1}}\max_{\begin{subarray}{c}\tau\in\Theta_{K^{-1}}\\ \measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\gtrsim K^{-1}\text{for all $1\leq i\leq A$}\end{subarray}}\int_{B_{K^{2}}^{n}(y)}|T^{\lambda}f_{\tau}|^{p}$
Given $U\subset B_{R}^{n}(0)$, we define the broad norm over $U$ as follows
$\|T^{\lambda}f\|_{\textup{BL}^{p}_{A}(U)}^{p}:=\sum_{B_{K^{2}}^{n}\in\mathcal{B}}\frac{|B_{K^{2}}^{n}\cap U|}{|B_{K^{2}}^{n}|}\mu(B_{K^{2}}^{n})$
The broad norm has the following properties; cf. [6, Section 4] for proofs.
Lemma 3.19.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition. Let $f_{1},f_{2}:\mathbb{R}^{n-1}\mapsto\mathbb{C}$. Suppose $1\leq p<\infty$, $A_{1},A_{2}\geq 1$ and $A=A_{1}+A_{2}$, then
$\|T^{\lambda}(f_{1}+f_{2})\|_{\textup{BL}^{p}_{A}(U)}\lesssim\|T^{\lambda}f_{1}\|_{\textup{BL}^{p}_{A_{1}}(U)}+\|T^{\lambda}f_{2}\|_{\textup{BL}^{p}_{A_{2}}(U)}.$
Lemma 3.20.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition. Let $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$. Suppose $1\leq p,p_{1},p_{2}<\infty$, $0\leq\alpha_{1},\alpha_{2}\leq 1$ satisfy $\alpha_{1}+\alpha_{2}=1$ and
$\frac{1}{p}=\frac{\alpha_{1}}{p_{1}}+\frac{\alpha_{2}}{p_{2}}.$
Suppose also $A_{1},A_{2}\geq 1$ and $A=A_{1}+A_{2}$. Then,
$\|T^{\lambda}f\|_{\textup{BL}^{p}_{A}(U)}\lesssim\|T^{\lambda}f\|_{\textup{BL}^{p_{1}}_{A_{1}}(U)}^{\alpha_{1}}\|T^{\lambda}f\|_{\textup{BL}^{p_{2}}_{A_{2}}(U)}^{\alpha_{2}}.$
3.8. Refined Decoupling.
The refined decoupling inequality first appeared in [8] as a strengthening of the classical decoupling inequality. In [4, Appendix A], the following refined decoupling inequality was established:
Theorem 3.21 ([4], Theorem A.3).
Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Suppose $1\leq R\leq\lambda^{1-\varepsilon}$. Assume that (3.1), (3.2), (3.22), and (3.23) hold. Let $\mathbb{T}\subset\mathcal{T}$ and define
$g:=\sum_{T\in\mathbb{T}}f_{T}.$
Let $w_{B_{R}}$ be a bump function satisfying $w_{B_{R}}(x)=1$ for $x\in B_{R}^{n}(0)$ and ${\rm supp}w_{B_{R}}\subset B_{2R}^{n}(0)$. Suppose that $\|T^{\lambda}f_{T}\|_{L^{q}(w_{B_{R}})}$ are comparable for all $T\in\mathbb{T}$. Let $Y$ be a union of finitely overlapping $K^{2}$-balls intersecting $B_{R}^{n}(0)$, each of which intersects at most $M$ tubes in $\mathbb{T}$. Then, for all $2\leq q\leq\frac{2n}{n-2}$ and all $\varepsilon>0$, we have
$\|T^{\lambda}g\|_{L^{q}(Y)}^{q}\lesssim_{\varepsilon,\phi}R^{\varepsilon}M^{\frac{q}{2}-1}\sum_{T\in\mathbb{T}}\|T^{\lambda}f_{T}\|_{L^{q}(w_{B_{R}})}^{q}+{\rm RapDec}(R)\,\|f\|_{2}^{q}.$
This theorem plays an important role in the restricted projection problem (see [3]), the local smoothing problem (see [4]), and the present work.
4. Two-ends reduction
In this section, we introduce two algorithms that generate the two-ends and broad structures required for the subsequent analysis. These algorithms may be regarded as streamlined versions of those developed in [4].
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Suppose $f\in\mathcal{S}(\mathbb{R}^{n-1})$ is supported in $\mathbb{A}$. Let $p:=2+\frac{8}{3n-5}$. Assume (3.1), (3.2), (3.22) and (3.23). Cover [R,R][-R,R] a finitely overlapping family of intervals $I$ of length $\sim R/K_{\circ}$. Let $\Omega=\{\omega\}$ denote the associated collection of horizontal regions, where each $\omega$ is of the form
$\mathbb{R}^{n-1}\times I$
In order to formalize the two-ends and broad structures, we introduce the following notions from [4, Section 6].
Definition 4.1 (Shaded incidence triple).
[4, Definition 6.1] Let $\Omega$ be the set of horizontal regions introduced above. A shaded incidence triple (or simply triple) $({\bf B},{\bf T};\mathcal{G})$ is the following:
  1. (1)
    ${\bf B}=\{B\}$ is a set of $K^{2}$-balls intersecting $B_{R}^{n}(0)$;
  2. (2)
    ${\bf T}=\{T\}$ is a set of $R$-planks in $B_{R}^{n}(0)$;
  3. (3)
    $\mathcal{G}$ is the shading map that $\mathcal{G}:{\bf T}\to P(\Omega)$. Here $P(X)$ is the collection of subsets of $X$, which is also called the power set of $X$.
Provided two shading maps $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, we say $\mathcal{G}_{1}\subset\mathcal{G}_{2}$ iff $\mathcal{G}_{1}(T)\subset\mathcal{G}_{2}(T)$ for all $T\in{\bf T}$. Given a triple $({\bf B},{\bf T};\mathcal{G})$, we consider the following incidence count:
$\mathcal{I}({\bf B},{\bf T};\mathcal{G}):=\#\bigl\{(B,T)\in{\bf B}\times{\bf T}:B\cap 100T\neq\emptyset,\;B\subset\bigcup\mathcal{G}(T)\bigr\}.$
For each $B\in{\bf B}$, we further define
${\bf T}(B;\mathcal{G}):=\bigl\{T\in{\bf T}:B\cap 100T\neq\emptyset,\;B\subset\bigcup\mathcal{G}(T)\bigr\},$
and for each $T\in{\bf T}$, we define
${\bf B}(T;\mathcal{G}):=\bigl\{B\in\mathcal{B}:B\cap 100T\neq\emptyset,\;B\subset\bigcup\mathcal{G}(T)\bigr\}.$
Definition 4.2.
[4, Definition 7.2] Given a triple $({\bf B},{\bf T};\mathcal{G})$, we denote
$L^{p}({\bf B},{\bf T};\mathcal{G};f):=\sum_{B\in{\bf B}}\|\sum_{T\in{\bf T}(B;\mathcal{G})}T^{\lambda}f_{T}\|_{L^{p}(B)}^{p}.$
Definition 4.3.
[4, Definition 6.2] For an $R$-plank $T$, we use $\theta(T)\in\Theta_{R^{-1/2}}$ to denote its directional cap. Let $S$ be a union of $\theta\in\Theta_{R^{-1/2}}$ and let $\bf T$} be a set of $R$-planks. We define
${\bf T}[S]:=\{T\in{\bf T}:\theta(T)\subset S\}.$
Definition 4.4 (Broad incidence).
[4, Definition 6.3] Given a triple $({\bf B},{\bf T};\mathcal{G})$, define
$\mathcal{I}_{\textup{br}}({\bf B},{\bf T};\mathcal{G}):=\inf_{\tau\in\Theta_{K^{-1}}}\mathcal{I}({\bf B},{\bf T}[(4\tau)^{c}];\mathcal{G}).$
For a set of planks $\bf T$}, we define the broad cardinality as
$_{\textup{br}}{\bf T}:=\min_{\tau\in\Theta_{K^{-1}}}\#{\bf T}[(4\tau)^{c}].$
4.1. Algorithm 1
We now describe Algorithm 1, which provides a procedure for refining a given triple $({\bf B},{\bf T},\mathcal{G})$, where $\bf T$} represents the collection of wave packets and $\bf B$} represents the integration domain. The output of the algorithm is a refined triple together with several auxiliary parameters.
Input: Given an incidence triple $({\bf B},{\bf T};\mathcal{G})$ and a $\nu>0$ such that for $B\in{\bf B}$,
$\big\|\sum_{T\in{\bf T}(B,\mathcal{G})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}\sim\nu.$
Pigeonholing and pruning: By dyadic pigeonholing, we can find some $\mu\lesssim R^{O(1)}$ and ${\bf B}_{1}\subset{\bf B}$ with
(4.1) $_{\textup{br}}{\bf T}(B;\mathcal{G})\sim\mu\qquad\textup{for all $B\in{\bf B}_{1}$}$
and
(4.2) $L^{p}({\bf B},{\bf T};\mathcal{G};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{1},{\bf T};\mathcal{G};f)$
Given $\lambda\lesssim R^{O(1)}$ and $\beta\lesssim K_{\circ}$, define
(4.3) $\mathcal{G}_{\lambda}(T):=\{\omega\in\mathcal{G}(T):\#\{B\in{\bf B}_{1}(T;\mathcal{G}):B\subset\omega\}\sim\lambda\}$
(4.4) ${\bf T}_{\lambda,\beta}=\{T\in{\bf T}:\#\mathcal{G}_{\lambda}(T)\sim\beta\}$
Fix $B\in{\bf B}_{1}$. By pigeonholing, there exists $\lambda(B)$ and $\beta(B)$ such that
$\nu\sim\big\|\sum_{T\in{\bf T}(B;\mathcal{G})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}\leq(\log R)^{O(1)}\big\|\sum_{T\in{\bf T}_{\lambda,\beta}(B;\mathcal{G}_{\lambda(B)})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}$
By dyadic pigeonholing again, we can find $\lambda$, $\beta$, $\nu_{1}$ and ${\bf B}_{2}\subset{\bf B}_{1}$ such that
$\lambda(B)\sim\lambda,\qquad\beta(B)\sim\beta,\qquad\big\|\sum_{T\in{\bf T}_{\lambda,\beta}(B;\mathcal{G}_{\lambda})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}\sim\nu_{1}\qquad\text{for all $B\in{\bf B}_{2}$}$
and
(4.5) $L^{p}({\bf B}_{1},{\bf T};\mathcal{G};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{2},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)$
Finally, we cover $B_{R}^{n}(0)$ by finite-overlapping $R^{1/2}$-cubes. By pigeonholing, we can find a dyadic number $m$ and a set $\mathcal{Q}$} of $R^{1/2}$-cubes such that
$_{\textup{br}}\bigcup_{B\in{\bf B}_{2},B\cap Q\neq\emptyset}{\bf T}_{\lambda,\beta}(B;\mathcal{G}_{\lambda})\sim m\qquad\text{for all $Q\in{\mathcal{Q}}$}$
and
(4.6) $L^{p}({\bf B}_{2},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{3},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)$
where
${\bf B}_{3}:=\{B:B\in{\bf B}_{2},B\cap\big(\bigcup_{Q\in{\mathcal{Q}}}Q\big)\neq\emptyset\}$
Define
$l:=\sup_{T\in{\bf T}_{\lambda,\beta}}\#\{Q\in{\mathcal{Q}}:Q\cap T\neq\emptyset\}$
Now we end Algorithm 1.
Some quantitative relationships: From (4.1),
(4.7) $\mathcal{I}({\bf B}_{1},{\bf T};\mathcal{G})=\sum_{B\in{\bf B}_{1}}\#{\bf T}(B;\mathcal{G})\geq\#{\bf B}_{1}\mu.$
From (4.3) and (4.4),
(4.8) $\mathcal{I}({\bf B}_{1},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda})=\sum_{T\in{\bf T}_{\beta}}\#{\bf B}_{1}(T;\mathcal{G}_{\lambda})\lesssim\#{\bf T}_{\lambda,\beta}\cdot\beta\lambda$
By (4.2), (4.5) and (4.6),
(4.9) $L^{p}({\bf B},{\bf T};\mathcal{G};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{3},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)$
An application of the classical hairbrush argument yields
Lemma 4.5.
[4, Lemma 6.4] Suppose $\beta>100$. Then
${\bf B}_{1}\gtrsim(KK_{\circ})^{-O(1)}lm\lambda$
4.2. Algorithm 2
Algorithm 2 repeatedly applies Algorithm 1 to establish additional useful quantitative relationships. Let $q:=2+\frac{4}{n-2}$. By a pigeonholing argument and Lemma 3.6, there exists a family ${\mathcal{B}}=\{B\}$ of finitely overlapping $K^{2}$-balls such that
$\|\sum_{T\in\mathbb{T},T\cap B\neq\emptyset}T^{\lambda}f_{T}\|_{L^{p}(B)}$
are comparable for all $B\in{\mathcal{B}}$ and
$\|\sum_{T\in\mathbb{T}}T^{\lambda}f_{T}\|_{L^{p}(B_{R}^{n+1})}^{p}\leq(\log R)^{O(1)}\sum_{B\in{\mathcal{B}}}\|\sum_{T\in\mathbb{T},T\cap B\neq\emptyset}T^{\lambda}f_{T}\|_{L^{p}(B)}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Combine this and (5.14), we see that
(4.10) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\leq(\log R)^{O(1)}\sum_{B\in{\mathcal{B}}}\|\sum_{T\in\mathbb{T},T\cap B\neq\emptyset}T^{\lambda}f_{T}\|_{L^{p}(B)}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Set $\mathbb{T}^{(0)}=\mathcal{T}$, ${\mathcal{B}}^{(0)}={\mathcal{B}}$ and
[…]
So, (4.10) can be rewritten as
(4.11) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\leq(\log R)^{O(1)}L^{p}({\mathcal{B}}^{(0)},\mathbb{T}^{(0)};\mathcal{G}^{(0)};f)+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Next, we apply Algorithm 1 iteratively. At the $i$-th step, we feed $({\mathcal{B}}^{(i-1)},\mathbb{T}^{(i-1)}$ $\mathcal{G}^{(i-1)};f)$ into Algorithm 1, which then outputs the following:
  1. (1)
    Parameters: $\mu^{(i-1)},\lambda^{(i-1)},\beta^{(i-1)},m^{(i-1)},l^{(i-1)}$.
  2. (2)
    Two families of $K^{2}$-balls: ${\mathcal{B}}_{3}^{(i-1)}\subset{\mathcal{B}}_{1}^{(i-1)}\subset{\mathcal{B}}^{(i-1)}$.
  3. (3)
    A set of planks $\mathbb{T}_{\lambda^{(i-1)},\beta^{(i-1)}}^{(i-1)}\subset\mathbb{T}^{(i-1)}$, a shading map $\mathcal{G}_{\lambda^{(i-1)}}^{(i-1)}\subset\mathcal{G}^{(i-1)}$, and a set of $R^{1/2}$-cubes ${\mathcal{Q}}^{(i-1)}$.
We set ${\mathcal{B}}^{(i)}:={\mathcal{B}}_{3}^{(i-1)}$, $\mathbb{T}^{(i)}:=\mathbb{T}_{\lambda^{(i-1)},\beta^{(i-1)}}^{(i-1)}$ and $\mathcal{G}^{(i)}:=\mathcal{G}_{\lambda^{(i-1)}}^{(i-1)}$.
Let $\kappa=R^{\varepsilon^{500}}$. Clearly, for all $i$ we have
$\mu_{i+1}\leq\mu_{i}\lesssim R^{O(1)}$
$\mathcal{I}({\mathcal{B}}_{1}^{(i+1)},\mathbb{T}^{(i+1)};\mathcal{G}^{(i+1)})\leq\mathcal{I}({\mathcal{B}}_{1}^{(i)},\mathbb{T}^{(i)};\mathcal{G}^{(i)})\lesssim R^{O(1)}$
Hence, for some $N\lesssim_{\varepsilon}1$
(4.12) $\mu_{N}\leq\kappa\mu_{N+1}$
(4.13) $\mathcal{I}({\mathcal{B}}_{1}^{(N)},\mathbb{T}^{(N)};\mathcal{G}^{(N)})\leq\kappa\mathcal{I}({\mathcal{B}}_{1}^{(N+1)},\mathbb{T}^{(N+1)};\mathcal{G}^{(N+1)})$
We now state an important quantitative relationship that will play a key role in the subsequent analysis.
Lemma 4.6.
If $\beta^{(N)}>100$, then
$\mu^{(N)}\lesssim R^{\varepsilon^{2}}\big(\#\mathbb{T}^{(N+1)}/l^{(N)}\big)^{1/2}$
Proof.
By Lemma 4.5, (4.7), (4.8) and (4.13)
(4.14) $\mu^{(N)}\leq(KK_{\circ})^{O(1)}\kappa\#\mathbb{T}^{(N+1)}\beta^{(N)}\big(m^{(N)}l^{(N)}\big)^{-1}$
By (4.12),
(4.15) $\mu^{(N)}\leq\kappa\mu^{(N+1)}\leq\kappa m^{(N)}$
Thus, the proof is completed by taking the geometric mean of (4.14) and (4.15). ∎
With a slight abuse of notation, we suppress the superscript and denote $\mu^{(N)}$, $\lambda^{(N)}$, $\beta^{(N)}$, $m^{(N)}$, $l^{(N)}$, ${\mathcal{B}}^{(N+1)}$, $\mathbb{T}^{(N+1)}$, $\mathcal{G}^{(N+1)}$ and ${\mathcal{Q}}^{(N)}$ simply by $\mu$, $\lambda$, $\beta$, $m$, $l$, $\mathcal{B}$}, $\mathbb{T}$, $\mathcal{G}$ and $\mathcal{Q}$}.
Since $N\lesssim_{\varepsilon}1$, (4.11) and (4.9) give
(4.16) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\leq(\log R)^{O_{\varepsilon}(1)}L^{p}({\mathcal{B}},\mathbb{T};\mathcal{G};f)+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
This finishes the two-ends redution.
5. Complete the proof
In this section, we complete the proof of Proposition 3.18. For the reader’s convenience, we restate the proposition below.
Proposition 5.1.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), (3.2), (3.22) and (3.23), then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{p}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{p}}$
$\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}:=\left(\frac{1}{|3\tilde{\theta}|}\int_{3\tilde{\theta}}|f|^{2}\right)^{1/2}$
holds uniformly for $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$, $1\leq R\leq\lambda^{1-\varepsilon}$ and $\varepsilon>0$.
We work under the same setting as in Section 4. Cover $B_{R}^{n}(0)$ by finite-overlapping $R/K_{\circ}$-balls $B_{k}$. For each $B_{k}$ define $\mathbb{T}_{k}$ by
$\mathbb{T}_{k}:=\{T\in\mathbb{T},100T\cap B_{k}\cap\big(\cup\mathcal{G}(T)\big)\neq\emptyset\}$
and $f_{k}$ by
$f_{k}:=\sum_{T\in\mathbb{T}_{k}}f_{T}$
Now we can rewrite (4.16) as
(5.1) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim(\log R)^{O_{\varepsilon}(1)}\sum_{B_{k}}\|T^{\lambda}f_{k}\|_{L^{p}(B_{k})}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
In the rest of this section, we will consider the following two cases separately:
  1. (1)
    One-end: $\beta\leq 100$.
  2. (2)
    Two-ends: $\beta>100$.
5.1. Case: One-end
In this subsection, we prove Proposition 3.18 under the assumption that $\beta\leq 100$. Applying the induction hypothesis to each $B_{k}$, we obtain
(5.2) $\|T^{\lambda}f_{k}\|_{L^{p}(B_{k})}^{p}\leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{K_{\circ}^{1/2}R^{-1/2}}}\|f_{k}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p}\\ \leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
By Lemma 3.16,
(5.3) $\max_{\theta\in\Theta_{R^{-1/2}}}\|f_{k}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\lesssim\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p-2}$
Combine this with (5.2),
$\|T^{\lambda}f_{k}\|_{L^{p}(B_{k})}^{p}\leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
By (5.1),
(5.4) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim C_{\varepsilon}K_{\circ}^{-p\varepsilon}(\log R)^{O_{\varepsilon}(1)}R^{p\varepsilon}\sum_{B_{k}}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Since $\beta\leq 100$, each wave packet $f_{T}$ appears in the sum defining $O(1)$ different functions $f_{k}$. Consequently, Lemma 3.7 yields
$\sum_{k}\|f_{k}\|_{2}^{2}\lesssim\|f\|_{2}^{2}$
Combine this with (5.4),
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim C_{\varepsilon}K_{\circ}^{-p\varepsilon}(\log R)^{O_{\varepsilon}(1)}R^{p\varepsilon}\|f\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Since $C_{\varepsilon}K_{\circ}^{-p\varepsilon}(\log R)^{O_{\varepsilon}(1)}$ can be arbitrarily small, the induction closes.
5.2. Two-ends: Reduction to Broad Estimate
Suppose $\beta>100$. Now we reduce Proposition 3.18 to broad case. By pigeonholing, we can find a $B_{k}$ such that
(5.5) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lessapprox K_{\circ}^{O(1)}\sum_{B\in{\bf B},B\subset B_{k}}\|T^{\lambda}f_{k}\|_{L^{p}(B)}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Let $A=10000$. For any $B_{K^{2}}^{n}(y)\in\mathcal{B}$ with $B_{K^{2}}^{n}(y)\cap B_{k}\neq\emptyset$, fix a choice of $v_{1},\dots,v_{A}\in S^{n-1}$ so that the minimum
$\min_{v_{1},\dots,v_{A}\in S^{n-1}}\max_{\begin{subarray}{c}\tau\in\Theta_{K^{-1}}\\ \measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\gtrsim K^{-2}\text{for all $1\leq i\leq A$}\end{subarray}}\int_{B_{K^{2}}^{n}(y)}|T^{\lambda}f_{k,\tau}|^{p}\\ \text{where}\quad f_{k,\tau}:=(f_{k})_{\tau}$
is attained. By Hölder’s inequality,
$\|T^{\lambda}f_{k}\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\lesssim\sum_{1\leq i\leq A}\|T^{\lambda}(\sum_{{\begin{subarray}{c}\tau\in\Theta_{K^{-1}}\\ \measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\lesssim K^{-2}\end{subarray}}}f_{k,\tau})\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\\ +K^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(B_{K^{2}}^{n}(y))}^{p}$
For each $i$, there are $O(1)$ $K^{-1}$-caps satisfies $\measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\lesssim K^{-2}$, so
$\|T^{\lambda}f_{k}\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\lesssim\sum_{\tau\in\Theta_{K^{-1}}}\|T^{\lambda}f_{k,\tau}\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\\ +K^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(B_{K^{2}}^{n}(y))}^{p}$
Let $X:=\bigcup_{B\in\mathcal{B},B\cap B_{k}\neq\emptyset}B$. By (5.5),
(5.6) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lessapprox K_{\circ}^{O(1)}\sum_{\tau\in\Theta_{K^{-1}}}\|T^{\lambda}f_{k,\tau}\|_{L^{p}(B_{R}^{n}(0))}^{p}\\ +\big(K_{\circ}K\big)^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Apply induction on scale, Lemma 3.17 and (5.3) to each $\|T^{\lambda}f_{k,\tau}\|_{L^{p}(B_{R}^{n}(0))}^{p}$,
$\|T^{\lambda}f_{k,\tau}\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}K^{\,2-(n-2)(p-2)-p\varepsilon}R^{p\varepsilon}\,\|f_{k,\tau}\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Combine this with (5.6),
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lessapprox_{\phi}K_{\circ}^{O(1)}K^{\,2-(n-2)(p-2)-p\varepsilon}R^{p\varepsilon}\,\|f\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}\\ +\big(K_{\circ}K\big)^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}\vskip 12.0pt plus 4.0pt minus 4.0pt$
If $p>2+\frac{2}{n-2}$, then the narrow term vanishes, and hence it remains to show:
(5.7) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}^{p}\lesssim_{\phi,\varepsilon}R^{\varepsilon}\|f\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}$
5.3. Two-ends: Broad Estimate
In this subsection we prove (5.7). Let $\alpha:=\frac{n-1}{3n-1}$ and thus
$\frac{1}{p}=\frac{\alpha}{2}+\frac{1-\alpha}{q}$
By Lemma 3.20, we have
(5.8) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}\lesssim\|T^{\lambda}f_{k}\|^{\alpha}_{\textup{BL}^{2}_{A/2}(X)}\|T^{\lambda}f_{k}\|^{1-\alpha}_{\textup{BL}^{q}_{A/2}(X)}$
Hence, it remains to estimate $\|T^{\lambda}f_{k}\|_{\textup{BL}^{2}_{A/2}(X)}$ and $\|T^{\lambda}f_{k}\|_{\textup{BL}^{q}_{A/2}(X)}$.
5.3.1. $L^{q}$-estimates using refined decoupling
Recall (4.1), for each $K^{2}$-ball $B$ intersecting $X$,
$_{\textup{br}}\{T\in\mathbb{T}_{k}:100T\cap B\neq\emptyset\}\leq\#_{\textup{br}}{\bf T}(B;\mathcal{G})\lesssim\mu.$
Hence, one can find $\tau^{\prime}\in\Theta_{K^{-1}}$ with
(5.9) $\{T\in\mathbb{T}_{k}:100T\cap B\neq\emptyset,\,\theta(T)\not\subset 4\tau^{\prime}\}\lesssim\mu$
By the definition of broad norm (3.24), there exists a $\tau(B)\not\subset 10\tau^{\prime}$ so that
(5.10) $\|T^{\lambda}f_{k}\|^{q}_{\textup{BL}^{q}_{A/2}(B)}\lesssim\|T^{\lambda}f_{k,\tau(B)}\|^{q}_{L^{q}(B)}$
Since $\tau(B)\not\subset 10\tau^{\prime}$, (5.9) gives
(5.11) $\{T\in\mathbb{T}_{k}:100T\cap B\neq\emptyset,\,3\theta(T)\cap 2\tau(B)\neq\emptyset\}\lesssim\mu.$
By pigeonholing and (5.10), there exists $\tau\in\Theta_{K^{-1}}$ and a union of $K^{2}$-balls $X^{\prime}\subset X$ such that
(5.12) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{q}_{A/2}(X)}^{q}\lesssim K^{O(1)}\|T^{\lambda}f_{k,\tau}\|_{L^{q}(X^{\prime})}^{q}$
Define $\bf T$} by
${\bf T}:=\{T\in\mathbb{T}_{k}:T\cap X^{\prime}\neq\emptyset,\,3\theta(T)\cap 2\tau\neq\emptyset\}$
Let $w_{B_{R}}$ be a bump function satisfying $w_{B_{R}}(x)=1$ for $x\in B_{R}^{n}(0)$ and ${\rm supp}w_{B_{R}}\subset B_{2R}^{n}(0)$. By pigeonholing, there exist dyadic numbers $u_{1}$$u_{2}$ and ${\bf T}^{\prime}\subset{\bf T}$ such that
(5.13) $\|T^{\lambda}{f_{k,\tau}}_{T}\|_{L^{q}(w_{B_{R}})}\sim u_{1},\qquad\|{f_{k,\tau}}_{T}\|_{L^{2}(\mathbb{R}^{n-1})}\sim u_{2},\qquad\text{for all $T\in{\bf T}^{\prime}$}.$
and
(5.14) $\|T^{\lambda}f_{k,\tau}\|_{L^{q}(X^{\prime})}^{q}\leq(\log R)^{O(1)}\|\sum_{T\in{\bf T}^{\prime}}T^{\lambda}{f_{k,\tau}}_{T}\|_{L^{q}(X^{\prime})}^{q}+{\rm RapDec}(R)\,\|f\|_{2}^{q},$
By (5.11), Theorem 3.21, Corollary 3.11 and (5.13),
(5.15) $\|\sum_{T\in{\bf T}^{\prime}}T^{\lambda}{f_{k,\tau}}_{T}\|_{L^{q}(X^{\prime})}^{q}\lessapprox_{\phi}\mu^{\frac{q}{2}-1}\#\mathbb{T}u_{2}^{q}+{\rm RapDec}(R)\,\|f\|_{2}^{q}$
By Lemma 3.15,
$u_{2}^{2}R^{\frac{n-2}{4}}|\mathbb{{\bf T}^{\prime}}|^{1/2}\lesssim\max_{\theta\in\Theta_{R^{-1/2}}}\|f_{k,\tau}\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}$
Combine this with (5.15),(5.14), (5.12) and Lemma 3.16,
(5.16) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{q}_{A/2}(X)}^{q}\lessapprox_{\phi}R^{-1/2}\mu^{\frac{2}{n-2}}\big(\#\mathbb{T}\big)^{-\frac{1}{n-2}}\|f\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,q-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{q}$
5.3.2. A simple $L^{2}$-estimate
By pigeonholing, there exists $\tau\in\Theta_{K^{-1}}$ and a union of $K^{2}$-balls $X^{\prime}\subset X$ such that
(5.17) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{2}_{A/2}(X)}^{2}\lesssim K^{O(1)}\|T^{\lambda}f_{k,\tau}\|_{L^{2}(X^{\prime})}^{2}\lesssim K^{O(1)}\sum_{Q\in{\mathcal{Q}}}\|T^{\lambda}f_{k,\tau}\|_{L^{2}(Q)}^{2}\\ =K^{O(1)}\sum_{Q\in{\mathcal{Q}}}\|T^{\lambda}\big(\sum_{T\in\mathbb{T},T\cap Q\neq\emptyset}{f_{k,\tau}}_{T}\big)\|_{L^{2}(Q)}^{2}+{\rm RapDec}(R)\,\|f\|_{2}^{2}$
For each $Q$, Corollary 3.12 and Lemma 3.7 give
$\|T^{\lambda}\big(\sum_{T\in\mathbb{T},T\cap Q\neq\emptyset}{f_{k,\tau}}_{T}\big)\|_{L^{2}(Q)}^{2}\lesssim_{\phi}R^{1/2}\sum_{T\in\mathbb{T},T\cap Q\neq\emptyset}\|{f_{k,\tau}}_{T}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}$
Combine this with (5.17),
$\|T^{\lambda}f_{k}\|_{\textup{BL}^{2}_{A/2}(X)}^{2}\lessapprox_{\phi}R^{1/2}\sum_{Q\in{\mathcal{Q}}}\sum_{T\in\mathbb{T},T\cap Q\neq\emptyset}\|{f_{k,\tau}}_{T}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}+{\rm RapDec}(R)\,\|f\|_{2}^{2}$
Since each $T\in\mathbb{T}$ intersects $O(l)$ cubes $Q$s, Lemma 3.7 yields
(5.18) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{2}_{A/2}(X)}^{2}\lessapprox_{\phi}lR^{1/2}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{2}+{\rm RapDec}(R)\,\|f\|_{2}^{2}$
5.3.3. Interpolation
We conclude the proof via an interpolation argument. By (5.8), (5.16) and (5.18),
(5.19) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}\lessapprox_{\phi}R^{-\frac{n-3}{12n-4}}l^{\frac{n-3}{6n-2}}\big(\mu\big(\#\mathbb{T}\big)^{-1/2}l^{1/2}\big)^{\frac{2}{3n-1}}+{\rm RapDec}(R)\,\|f\|_{2}^{2}$
By Lemma 4.6,
(5.20) $\mu\big(\#\mathbb{T}\big)^{-1/2}l^{1/2}\lesssim R^{\varepsilon^{2}}$
Since $l\lesssim R^{1/2}$, (5.7) follows by combining (5.19) and (5.20).
6. An $\varepsilon$-removal lemma
In this section, we show that Proposition 1.2 implies Theorem 1.1. The $R^{\varepsilon}$–loss appearing in the linear estimates of Proposition 1.2 can be eliminated away from the endpoint by invoking an $\varepsilon$-removal lemma of the type introduced in [12]. Although the precise form of the required lemma does not seem to appear explicitly in the literature, it can be obtained through a minor adaptation of the arguments in [12] and [7].
Suppose $p_{0}\geq 2$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition. We are given that the estimate
(6.1) $\|T^{\lambda}f\|_{L^{p}(B^{n}_{R})}\lesssim_{\phi,a,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
holds uniformly for $1\leq R\leq\lambda$ whenever $p\geq p_{0}$. Under this assumption, we aim to show that the global estimate
(6.2) $\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{\phi,a,p}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
holds for all $p>p_{0}$.
We begin by reducing (6.2) to estimates over sparse collections of $R$-balls.
Definition 6.1.
[12] Let $R\geq 1$. A collection $\{B(x_{j},R)\}_{j=1}^{N}$ of $R$-balls in $\mathbb{R}^{n}$ is sparse if the centres $\{x_{1},\dots,x_{N}\}$ are $(RN)^{\bar{C}}$-separated. Here $\bar{C}\geq 1$ is a sufficiently large absolute constant.
Lemma 6.2.
[7, Lemma 12.2] Suppose $p_{0}\geq 2$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition. Assume that (6.1) holds uniformly for $1\leq R\leq\lambda$ whenever $p\geq p_{0}$. To prove (6.2) for all $p>p_{0}$ it suffices to show that for all $\varepsilon>0$ the estimate
(6.3) $\|T^{\lambda}f\|_{L^{p_{0}}(S)}\lesssim_{\varepsilon,\phi,a}R^{\varepsilon}\|f\|_{L^{p_{0}}(B^{n-1})}$
holds whenever $R\geq 1$ and $S\subseteq\mathbb{R}^{n}$ is a union of $R$-balls belonging to a sparse collection.
The main result of this subsection is stated below and can be deduced by a minor modification of the argument in [7, Lemma 12.3].
Proposition 6.3.
Suppose $p_{0}\geq 2$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition. Assume that (6.1) holds uniformly for $1\leq R\leq\lambda$ whenever $p\geq p_{0}$. Then for all $\varepsilon>0$ the estimate
(6.4) $\|T^{\lambda}f\|_{L^{p_{0}}(S)}\lesssim_{\varepsilon,\phi,a}R^{\varepsilon}\|f\|_{L^{p_{0}}(B^{n-1})}$
holds whenever $R\geq 1$ and $S\subseteq\mathbb{R}^{n}$ is a union of $R$-balls belonging to a sparse collection.
Proof.
Let $\{B(x_{j},R)\}_{j=1}^{N}$ be a sparse collection of balls whose union equals the set $S$. It is clear that we may assume $R\ll\lambda$ and that each ball $B(x_{j},R)$ intersects the $x$-support of $a^{\lambda}$. Moreover, let $c_{\mathrm{diam}}>0$ be a sufficiently small constant, chosen to meet the requirements of the argument below. By applying a partition of unity, we may further assume that $\operatorname{diam}X<c_{\mathrm{diam}}$, and hence
(6.5) $\frac{|x_{j_{1}}-x_{j_{2}}|}{\lambda}\lesssim c_{\mathrm{diam}}\qquad\textrm{for all $1\leq j_{1},j_{2}\leq N$.}$
Fix a function $\eta\in C^{\infty}(\mathbb{R}^{n-1})$ such that $0\leq\eta\leq 1$, ${\rm supp}\eta\subset B^{n-1}(0,1)$, and $\eta(z)=1$ for all $z\in B^{n-1}(0,1/2)$. Let $R_{1}:=CR$, where $C\geq 1$ is a sufficiently large constant, and define $\eta_{R_{1}}(z):=\eta(z/R_{1})$. Let $R_{2}:=(RN)^{\bar{C}}$. In addition, let $\psi\in C_{c}^{\infty}(\mathbb{R}^{n-1})$ satisfy $0\leq\psi\leq 1$, ${\rm supp}\psi\subset\Omega$, and $\psi(\omega)=1$ for all $\omega$ in the $\omega$-support of $a^{\lambda}$.
Fix $1\leq j\leq N$. We decompose
$e^{2\pi i\phi^{\lambda}(x_{j};\,\cdot\,)}\psi f=P_{j}f+\bigl(e^{2\pi i\phi^{\lambda}(x_{j};\,\cdot\,)}\psi f-P_{j}f\bigr)=:P_{j}f+f_{j,\infty},$
where
$P_{j}f:=\widehat{\eta}_{R_{1}}*\bigl[e^{2\pi i\phi^{\lambda}(x_{j};\,\cdot\,)}\psi f\bigr].$
As shown in [7, Proof of Lemma 12.3], it suffices to show that the inequality
(6.6) $\|\sum_{j=1}^{N}P_{j}^{*}g_{j}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}\lesssim\sum_{j=1}^{N}\|g_{j}\|_{L^{2}(\mathbb{R}^{n-1})}^{2}.$
holds uniformly for all $g_{1},g_{2},\dots,g_{N}\in L^{2}(\mathbb{R}^{n-1})$. The left-hand side of the above expression equals
(6.7) $\sum_{j_{1},j_{2}=1}^{N}\int_{\mathbb{R}^{n-1}}\overline{G_{j_{1},j_{2}}(\omega)}\hat{\eta}_{R_{1}}\ast g_{j_{1}}(\omega)\overline{\hat{\eta}_{R_{1}}\ast g_{j_{2}}(\omega)}\ d\omega$
where
$G_{j_{1},j_{2}}(\omega):=e^{2\pi i(\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega))}\psi(\omega)^{2}.$
By Plancherel’s theorem, each summand in (6.7) can be written as
$\int_{\mathbb{R}^{n-1}}\overline{\check{G}_{j_{1},j_{2}}(z)}(\eta_{R_{1}}\check{g}_{j_{1}})\ast(\eta_{R_{1}}\check{g}_{j_{2}})^{\sim}(z)dz;$
here $(\eta_{R_{1}}\check{g}_{j_{2}})^{\sim}(z):=\overline{(\eta_{R_{1}}\check{g}_{j_{2}})(-z)}$.
Fix $1\leq j_{1},j_{2}\leq N$ with $j_{1}\neq j_{2}$, and choose $z\in\mathbb{R}^{n-1}$ satisfying $|z|\lesssim R_{1}\ll R_{2}\leq|x_{j_{2}}-x_{j_{1}}|$. Consider
$\check{G}_{j_{1},j_{2}}(z)=\int_{\mathbb{R}^{n-1}}e^{2\pi i(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega))}\psi(\omega)^{2}d\omega.$
Next, we decompose $B^{n-1}_{1}(0)$ into $O_{\phi}(1)$ finitely overlapping open sets $D$ such that
(6.8) $\int_{\mathbb{R}^{n-1}}e^{2\pi i\bigl(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega)\bigr)}\,\psi(\omega)^{2}\,\psi_{D}(\omega)d\omega\;\lesssim\;R_{2}^{-\frac{n-2}{2}},$
where $\{\psi_{D}\}$ is a smooth partition of unity subordinate to this covering. Once we have such a decomposition, (6.6) follows by an application of the Hausdorff–Young inequality together with Hölder’s inequality.
We now start the partitioning. Let $C^{\prime}>0$ be a large constant and
$D_{0}:=\{\omega:\omega\in B_{n-1}^{1}(0),\big|\pm\frac{x_{j_{2}}-x_{j_{1}}}{|x_{j_{2}}-x_{j_{1}}|}-G^{\lambda}(x_{j_{1}};\omega)\big|>C^{\prime}c_{\mathrm{crit}}\}$
$U_{0}:=\{\omega:\omega\in B_{n-1}^{1}(0),\big|\pm\frac{x_{j_{2}}-x_{j_{1}}}{|x_{j_{2}}-x_{j_{1}}|}-G^{\lambda}(x_{j_{1}};\omega)\big|<2C^{\prime}c_{\mathrm{crit}}\}$
Define
$M(\omega_{0}):=\partial_{\omega\omega}\langle\partial_{x}\phi^{\lambda}(x_{j_{1}};\omega),G^{\lambda}(x;\omega_{0})\rangle|_{\omega=\omega_{0}}$
and let $M_{i}(\omega_{0})$ denote the matrix obtained by deleting the $\omega_{i}$-row and column of $M(\omega_{0})$. Let $c^{\prime}>0$ be a sufficiently small constant. Define
(6.9) $U_{i}:=\{\omega:\omega\in B_{n-1}^{1}(0),\det M_{i}(\omega)>c^{\prime}\}$
$D_{i}:=U_{i}\cap U_{0}$
The sets $D_{0},D_{1},\dots,D_{n-1}$ form an open cover of $B_{n-1}^{1}(0)$. Let $\{\psi_{D_{0}},\psi_{D_{1}},\dots,\psi_{D_{n-1}}\}$ be a smooth partition of unity subordinate to this covering.
We now verify (6.8). If $\omega\in D_{0}$, then $x_{j_{1}}-x_{j_{2}}$ makes a quantitative angle with the kernel of $\phi^{\lambda}(x_{j_{1}};\omega)$, and hence
$\partial_{\omega}\bigl(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega)\bigr)=\langle\partial_{\omega x}\phi^{\lambda}(x_{j_{1}};\omega),x_{j_{1}}-x_{j_{2}}\rangle\\ +O(c_{\mathrm{diam}}|x_{j_{1}}-x_{j_{2}}|)\gtrsim|x_{j_{1}}-x_{j_{2}}|$
Then the method of stationary phase gives
$\int_{\mathbb{R}^{n-1}}e^{2\pi i\bigl(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega)\bigr)}\,\psi(\omega)^{2}\,\psi_{D_{0}}(\omega)d\omega\;\lesssim\;|x_{j_{1}}-x_{j_{2}}|^{-(n-1)}\lesssim R_{2}^{-(n-1)}$
If $\omega\in D_{i}$ for some $i\in[1,n-1]$, then we have
$\partial_{\omega^{\prime}\omega^{\prime}}\bigl(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega)\bigr)=|x_{j_{1}}-x_{j_{2}}|M_{i}(\omega)+O(C^{\prime}c_{\mathrm{crit}}|x_{j_{1}}-x_{j_{2}}|)$
where $\omega^{\prime}=(\omega_{1},\dots,\omega_{i-1},\omega_{i+1},\dots,\omega_{n-1})$. Recall (6.9), if $C^{\prime}c_{\mathrm{crit}}$ is small enough, then
$\det\partial_{\omega^{\prime}\omega^{\prime}}\bigl(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega)\bigr)\gtrsim|x_{j_{1}}-x_{j_{2}}|^{n-2}$
By higher dimensional versions of van der Corput’s lemma,
$\int_{\mathbb{R}^{n-1}}e^{2\pi i\bigl(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega)\bigr)}\,\psi(\omega)^{2}\,\psi_{D_{i}}(\omega)d\omega\\ =\int_{1}^{1}\int_{\mathbb{R}^{n-2}}e^{2\pi i\bigl(\langle z,\omega\rangle+\phi^{\lambda}(x_{j_{1}};\omega)-\phi^{\lambda}(x_{j_{2}};\omega)\bigr)}\,\psi(\omega)^{2}\,\psi_{D_{i}}(\omega)d\omega^{\prime}d\omega_{i}\\ \lesssim|x_{j_{1}}-x_{j_{2}}|^{-\frac{n-2}{2}}\lesssim R_{2}^{-\frac{n-2}{2}}$